# A 3.3904-Competitive Online Algorithm for List Update with Uniform Costs

### v1

@@ Abstract
We consider the List Update problem where the cost of each swap is assumed to be 1. This is in contrast to the “standard” model, in which an algorithm is allowed to swap the requested item with previous items for free. We construct an online algorithm Full-Or-Partial-Move (Fpm), whose competitive ratio is $R\approx 3.3904$, improving over the previous best known bound of $4$.
@@
List update, online algorithms, competitive analysis
@@ 1 Introduction
In the online List Update problem, the objective is to maintain a set of items stored in a linear list in response to a sequence of access requests. The cost of accessing a requested item is equal to its distance from the front of the list. After each request, an algorithm is allowed to rearrange the list by executing any number of swaps of adjacent items. In the model introduced by Sleator and Tarjan in their seminal 1985 paper on competitive analysis [25], an algorithm can repeatedly swap the requested item with its preceding item at no cost. These swaps are called free. All other swaps are called paid and have cost $1$ each. As in other problems involving self-organizing data structures (see [7]), the objective of an algorithm is to judiciously adjust the list at each step to assure that, independently of future requests, its total cost is never too big in comparison to the optimum cost. If an algorithm’s cost is always within a factor $c$ of the optimum, it is said to be $c$-competitive, and this constant $c$ is referred to as its competitive ratio.
Sleator and Tarjan proved that Algorithm Move-to-Front (Mtf), that after each request moves the requested item to the front of the list, is $2$-competitive. This ratio is known to be optimal if the number of items is unbounded. Their work culminated prior extensive studies of List Update, that included experimental results, probabilistic approaches, and earlier attempts at amortized analysis (see [14] and the references therein).
As shown in subsequent work, Mtf is not unique — there are other strategies that achieve ratio $2$. One example is algorithm TimeStamp from [2]. Anderson et al. [9] showed that there is a large class of $2$-competitive algorithms based on work functions. In fact, it is known that there are infinitely many algorithms that achieve ratio $2$ [12, 21].
@@
Following [24], we will refer to the cost model of [25] as standard, and we will denote it here by $\textsf{LUP}_{0}$. This model has been questioned in the literature for not accurately reflecting true costs in some implementations [24, 23, 22, 18], with the concept of free swaps being one of the main concerns.
@@
A natural approach to address this concern, considered in some later studies of the List Update problem, is to simply charge cost $1$ for any swap. (See, for example, [24, 4, 11].) We will call it here the uniform cost model and denote it $\textsf{LUP}_{1}$. Changing the cost of free swaps from 00 to $1$ at most doubles any algorithm’s cost, so Mtf is no worse than $4$-competitive for $\textsf{LUP}_{1}$, and it is also easy to show that this bound is tight. A general lower bound of $3$ on the competitive ratio of deterministic algorithms for $\textsf{LUP}_{1}$ was given by Reingold et al. [24]. Natural adaptations of $2$-competitive algorithms for $\textsf{LUP}_{0}$ can be shown to have ratio larger than $3$ (see the discussion in section 5), and no algorithm is known to beat ratio $4$. In summary, in contrast to the standard model, closing or reducing this $3$-to-$4$ gap between the lower and upper bound for $\textsf{LUP}_{1}$ remains elusive [^1] The authors of [21] claimed an upper bound of $3$ for $\textsf{LUP}_{1}$, but discovered later that their proof was not correct (personal communication with S. Kamali)..
Our results.
Addressing this open problem, we develop an online algorithm for $\textsf{LUP}_{1}$ with competitive ratio ${\frac{1}{8}}\!\cdot\!(23+\sqrt{17})\approx 3.3904$, significantly improving the previous upper bound of $4$. Our algorithm, called Fpm, is based on a combination of several ideas. One involves the movement of the requested item. It can be shown (see section 5) that any algorithm that chooses either not to move at all or to move to front cannot be $3$-competitive. For this reason, Fpm will sometimes make a partial move to a suitably chosen target location in the list. Another idea involves keeping track, for each pair of items, of the work function for the two-item subsequence consisting of these items. These work functions are used in two ways. On one hand, they indicate, roughly, which relative order between the items in each pair is more likely in an optimal solution. The algorithm uses this information to determine the target location of each item. On the other hand, the simple sum of all these pair-based work functions is a lower bound on the optimum cost, which comes in handy in the competitive analysis of Fpm.
@@
Fpm remains $3.3904$-competitive even for the partial cost function, where the cost to access location $\ell$ is $\ell-1$, instead of the full cost of $\ell$ used in the original definition of List Update in [25]. Both functions have been alternatively used in the literature, depending on context and convenience. For any online algorithm, its competitive ratio for partial costs is at least as large as for full costs, although the difference typically vanishes if the list size is unbounded. We present our analysis in terms of partial costs. Further, the $3.3904$-competitiveness proof for Fpm easily extends to the dynamic scenario of List Update that also allows operations of insertions and deletions as in the original definition from [25].
The most intriguing question left open in our work is whether competitive ratio $3$ can be achieved. We discuss it in section 5. As explained there, we verified computationally that $3$-competitive algorithms exist for lists with up to $6$ items. On the other hand, several natural strategies (including our Fpm) fail to produce a general $3$-competitive algorithm.
Other related work.
Randomization is a natural tool in the online setting, and for many online problems it can significantly improve the competitive ratio. For $\textsf{LUP}_{0}$, the study of randomized algorithms was initiated by Irani [19] who showed an upper bound of $\approx 1.94$, the first bound below $2$. This was subsequently improved in [24, 2, 6], with the last paper, by Albers et al., giving the best currently known upper bound of $1.6$. This is already quite close to the lower bound of $1.5$, given by Teia [26] in 1993. In fact, in the partial-cost model there is a matching lower bound [8] for so-called projective algorithms, whose computation is uniquely determined by their behavior on two-item instances. For the uniform cost model ($\textsf{LUP}_{1}$), Reingold et al. [24] show an upper bound of $\approx 2.64$. Recent work by Albers and Janke [4] implies that this ratio can be improved to $\approx 2.4$.
Other cost models have been investigated. In the earlier mentioned papers [24, 4] on randomized algorithms, the authors generalize the uniform cost function by distinguishing between the cost of $1$ for following a link during search and the cost of $d\geq 1$ for a swap. A completely different approach is discussed in [22, 18], where an algorithm is allowed to completely rearrange the list at the cost that is only linear with respect to the position of the accessed item. Yet another approach is discussed in [23]. In these last two models no algorithm can be constant-competitive.
@@
Speaking to the centrality of the List Update problem within the area of competitive analysis, a variety of its variants has been investigated in the literature over the last forty years, including models with lookahead [1], locality of reference [10, 5], parametrized approach [17] and algorithms with advice [15]. A particularly interesting model was proposed recently by Azar et al. [11], where an online algorithm is allowed to postpone serving some requests, but is either required to serve it by a specified deadline or has to pay some delay penalty.
There is also a number of survey papers on the List Update problem, where the reader can find more information about this problem and its variants [3, 20, 21].
@@ 2 Preliminaries
An algorithm has to maintain a list of items, while a sequence $\sigma$ of access requests is presented to an algorithm in an online manner. In each step $t\geq 1$ an algorithm is presented an access request $\sigma^{t}$ to an item in the list. If this item is in a location $\ell$, the algorithm incurs cost $\ell-1$ to access it. (The locations in the list are indexed $1,2,...$.) Afterwards, the algorithm may change the list configuration by performing an arbitrary number of swaps of neighboring items, each of cost $1$.
For any algorithm $A$, we denote its cost for processing a sequence $\sigma$ by $A(\sigma)$. Furthermore, we denote the optimal algorithm by Opt.
@@
For an input $\sigma$ and a pair $\{x,y\}\in\mathcal{P}$, $\sigma_{xy}$ is the subsequence of $\sigma$ restricted to requests to items $x$ and $y$ only. Whenever we say that an algorithm serves input $\sigma_{xy}$, we mean that an algorithm has to maintain a list of two items, $x$ and $y$.
@@ 2.1 Work Functions
Figure 1: The evolution of the work function $W^{xy}$. Initially, $W^{xy}$ has its minimum in the state $yx$ and $y\prec x$ (in the list of an online algorithm). Thus, the mode of the pair $\{x,y\}$ is $\alpha$. Next, because of the two requests to $x$, the value of $W^{xy}(yx)$ is incremented, while the value at $xy$ remains intact. The mode is thus changed from $\alpha$ to $\beta$ and then to $\gamma$. Finally, when an algorithm moves item $x$ before $y$, the mode of the pair $\{x,y\}$ changes to $\alpha$. The value of $w^{xy}$ (the average of $W^{xy}(xy)$ and $W^{xy}(yx)$) increases by ${\frac{1}{2}}$ whenever the mode changes due to a request.
@@ 2.2 Lower Bound on OPT
2.2 gives us a convenient tool to lower bound $\textsl{{Opt}}(\sigma)$. We define the cost of pair-based Opt in step $t$ as $\sum_{\{x,y\}\in\mathcal{P}}\Delta_{t}w^{xy}$. For a given request sequence $\sigma$, the sum of these costs over all steps when processing $\sigma$ is a lower-bounds on the actual value of $\textsl{{Opt}}(\sigma)$.
@@ 4.3 State Transitions
While we defined 12 possible states (3 modes $\times$ 4 flavors), we will show that $\alpha^{n}$, $\gamma^{d}$, and $\gamma^{o}$ never occur. This clearly holds at the very beginning, as all pairs are then in state $\alpha^{d}$.
For succinctness, we also combine some of the remaining states, reducing the number of states to the following six: $\alpha^{d}$, $\beta^{d}$, $\alpha^{oe}$, $\beta^{o}$, $\beta^{ne}$ and $\gamma^{ne}$. (For example, a pair is in state $\alpha^{oe}$ if it’s in state $\alpha^{o}$ or $\alpha^{e}$.) In the following we analyze the transitions between them. We start with a simple observation.
@@
The pair $\{z^{*},y\}$ is not swapped due to the request, and thus, by subsection 4.2, the mode transition are $\alpha\to\alpha$, $\beta\to\alpha$, $\gamma\to\beta$.
@@ 4.4 Amortized Analysis
For succinctness, wherever it does not lead to ambiguity, we omit subscripts $t$, i.e., write $\Delta\Phi_{xy}$ and $\Delta w^{xy}$ instead of $\Delta_{t}\Phi_{xy}$ and $\Delta_{t}w^{xy}$. We also omit superscripts $t$ in $P^{t}_{1}$, $P^{t}_{2}$, and $P^{t}_{3}$.
@@
  • $\max\{\beta^{d},\beta^{o}\}\leq\gamma^{ne}+\textstyle{\frac{1}{2}}R$,
@@ 4.4.4 Proof of $R$-competitiveness
Next, assume Fpm performs a full move. Then, $\Delta\textsl{{Fpm}}=[2,2,2,2,2,2]\odot V$. Again by Table 1,
@@
Recall that Fpm is defined exactly to choose the move that minimizes $\Delta\textsl{{Fpm}}+\sum_{y\prec z^{*}}\Delta\Phi_{z^{*}y}$.
@@
[…]
@@
[…]
@@
5 Discussion
The question whether there is a $3$-competitive algorithm for List Update with uniform costs remains open. Here we discuss some side results and observations related to this question. Unless otherwise explicitly stated, we assume the partial cost model in this section (that is, accessing location $\ell$ costs $\ell-1$).
While our analysis of Fpm gives an upper bound of $3.3904$, we do not know what is the best ratio achievable with its general approach. For the specific choice of parameters we use in the paper, we verified that Fpm is $3$-competitive for lists of length $3$ (see Appendix D), but not better than $3.04$-competitive for lists of length $5$ (see Appendix E).
For lists of length $3$ there is in fact a much simpler algorithm that uses only a small number of states (at most $6$) to keep track of the past history. Our computational experiments indicate that $3$-competitive algorithms exist for lists with up to $6$ items (see Appendix B).
On the other hand, we also have explored several natural adaptations of algorithms for $\textsf{LUP}_{0}$, showing that none of them has ratio $3$ for $\textsf{LUP}_{1}$. Starting with Algorithm Move-to-Front (Mtf), its $4$-competitiveness follows from its $2$-competitiveness for $\textsf{LUP}_{0}$. The matching lower bound of $4$ is also easy: repeatedly request the last item in the list. Ignoring additive constants, the algorithm pays twice the length of the list at each step, while the static algorithm’s average cost per step is only half of that length. (For partial cost this works even with two items. For full costs, one needs to use arbitrarily long lists.)
The intuition why Mtf performs poorly is that it moves the requested items to front “too quickly”. For the particular strategy in the last paragraph (that always requests the last item), ratio $3$ can be obtained by moving the items to front only every second time they are requested. This algorithm, called Dbit, is a deterministic variant of algorithm Bit from [24] and a special case of algorithm Counter in [24] (with $\ell=2$). Algorithm Dbit was seen as a candidate for optimality in the past (see Footnote 1 in section 1). However, in Appendix C we show, that Dbit is not better than $4$-competitive, and not better than $3.302$-competitive in the full cost model.
The strategies discussed above, including Mtf, Dbit and Counter, either leave the requested item at its current location or move it to front. Generalizing the above bounds, we can show that any strategy limited to such moves cannot achieve ratio better than $3.25$, even for just three items (see subsection C.1). A naive fix would be to, say, move the requested item always half-way towards front. This algorithm is even worse: its competitive ratio is at least $6$ (Appendix C). (Our algorithm Fpm overcomes this difficulty by a much more sophisticated choice of the target location for the requested item.)
Anderson et al. [9] show that the work-function algorithm is $2$-competitive in the $\textsf{LUP}_{0}$ model. Our computer-aided calculation of its performance for short lists indicates that its ratio is $3$ for up to $4$ items, but it is larger than $3$ for $5$ items (see Appendix B).
To shed more light on the $\textsf{LUP}_{1}$ problem, one promising research direction is to examine in depth the concept of projective algorithms, as defined in [8]. The computation of a projective algorithm is uniquely determined by its behavior on two-item instances. As a result, to be $c$-competitive such an algorithm needs to be simultaneously $c$-competitive against each two-item subsequence of the request sequence. This property greatly simplifies competitive analysis, eliminating the need to keep track of the optimum cost for the full instance. Our algorithm Fpm is not projective. We do not know what is the best competitive ratio for $\textsf{LUP}_{1}$ achievable with a projective algorithm.

### v3

@@ Abstract
We consider the List Update problem where the cost of each swap is assumed to be 1. This is in contrast to the “standard” model, in which an algorithm is allowed to swap the requested item with previous items for free. We construct an online algorithm Full-Or-Partial-Move (Fpm), whose competitive ratio is at most $3.3904$, improving over the previous best known bound of $4$.
@@
List update, work functions, amortized analysis, online algorithms, competitive analysis
@@ 1 Introduction
In the online List Update problem [3, 21, 22], the objective is to maintain a set of items stored in a linear list in response to a sequence of access requests. The cost of accessing a requested item is equal to its distance from the front of the list. After each request, an algorithm is allowed to rearrange the list by performing an arbitrary number of swaps of adjacent items. In the model introduced by Sleator and Tarjan in their seminal 1985 paper on competitive analysis [26], an algorithm can repeatedly swap the requested item with its preceding item at no cost. These swaps are called free. All other swaps are called paid and have cost $1$ each. As in other problems involving self-organizing data structures [7], the goal is to construct an online algorithm, i.e., operating without the knowledge of future requests. The cost of such an algorithm is compared to the cost of the optimal offline algorithm; the ratio of the two costs is called the competitive ratio and is subject to minimization.
Sleator and Tarjan proved that the algorithm Move-To-Front (Mtf), which after each request moves the requested item to the front of the list, is $2$-competitive [26]. This ratio is known to be optimal if the number of items is unbounded. Their work was the culmination of previous extensive studies of list updating, including experimental results, probabilistic approaches, and earlier attempts at amortized analysis (see [15] and the references therein).
As shown in subsequent work, Mtf is not unique — there are other strategies that achieve ratio $2$, such as TimeStamp [2] or algorithms based on work functions [9]. In fact, there are infinitely many algorithms that achieve ratio 22 [14, 22].
@@
Following [25], we will refer to the cost model of [26] as standard, and we will denote it here by $\textsf{LUP}_{\mathrm{S}}$. This model has been questioned in the literature for not accurately reflecting true costs in some implementations [25, 24, 23, 19], with the concept of free swaps being one of the main concerns.
@@
A natural approach to address this concern, considered in some later studies (see, e.g., [25, 7, 22, 4, 13]), is simply to charge cost $1$ for any swap. We will call it here the uniform cost model and denote it $\textsf{LUP}_{1}$. A general lower bound of $3$ on the competitive ratio of deterministic algorithms for $\textsf{LUP}_{1}$ was given by Reingold et al. [25]. Changing the cost of free swaps from 0 to $1$ at most doubles the cost of any algorithm, so Mtf is no worse than $4$-competitive for $\textsf{LUP}_{1}$. Surprisingly, no algorithm is known to beat Mtf, i.e., achieve ratio lower than 44.111The authors of [22] claimed an upper bound of $3$ for $\textsf{LUP}_{1}$, but later discovered that their proof was not correct (personal communication with S. Kamali)..
Our main result.
To address this open problem, we develop an online algorithm Full-Or-Partial-Move (Fpm) for $\textsf{LUP}_{1}$ with competitive ratio ${\frac{1}{8}}\!\cdot\!(23+\sqrt{17})\approx 3.3904$, significantly improving the previous upper bound of $4$.
@@
Our algorithm Fpm remains $3.3904$-competitive even for the partial cost function, where the cost of accessing location $\ell$ is $\ell-1$, instead of the full cost of $\ell$ used in the original definition of List Update [26]. Both functions have been used in the literature, depending on context and convenience. For any online algorithm, its partial-cost competitive ratio is at least as large as its full-cost ratio, although the difference typically vanishes when the list size is unbounded. We present our analysis in terms of partial costs.
Fpm remains $3.3904$-competitive also in the dynamic scenario of List Update that allows operations of insertions and deletions, as in the original definition in [26] (see Appendix A).
Technical challenges and new ideas.
The question whether ratio $3$ can be achieved remains open. We also do not know if there is a simple algorithm with ratio below $4$. We have considered some natural adaptions of Mtf and other algorithms that are $2$-competitive for $\textsf{LUP}_{\mathrm{S}}$, but for all we were able to show lower bounds higher than $3$ for $\textsf{LUP}_{1}$.
As earlier mentioned, Mtf is $4$-competitive for $\textsf{LUP}_{1}$. It is also easy to show that its ratio is not better than $4$: repeatedly request the last item in the list. Ignoring additive constants, the algorithm pays twice the length of the list at each step, while any algorithm that just keeps the list in a fixed order, pays only half the length on the average.
The intuition why Mtf performs poorly is that it moves the requested items to front “too quickly”. For the aforementioned adversarial strategy against Mtf, ratio $3$ can be obtained by moving the items to front only every other time they are requested. This algorithm, called Dbit, is a deterministic variant of algorithm Bit from [25] and a special case of algorithm Counter in [25], and it has been also considered in [22]. In Appendix B, we show that Dbit is not better than $4$-competitive in the partial cost model, and not better than $3.302$-competitive in the full cost model.
One can generalize these approaches by considering a more general class of algorithms that either leave the requested item at its current location or move it to the front. In Appendix B we show that such a strategy cannot achieve ratio better than $3.25$, even for just three items. A naive fix would be, for example, to always move the requested item half-way towards front. This algorithm is even worse: its competitive ratio is at least $6$ (see Appendix B).
We also show (see section 5) via a computer-aided argument that, for $\textsf{LUP}_{1}$, the work function algorithm’s competitive ratio is larger than $3$, even for lists of length $5$. This is in contrast to the its performance for the standard model, where it achieves optimal ratio $2$ [9].
Our algorithm Fpm overcomes the difficulties mentioned above by combining a few new ideas. The first idea is a more sophisticated choice of the target location for the requested item. That is, aside from full moves that move the requested items to the list front, Fpm sometimes performs a partial move to a suitably chosen target location in the list. This location roughly corresponds to the front of the list when this item was requested earlier.
The second idea is to keep track, for each pair of items, of the work function for the two-item subsequence consisting of these items. These work functions are used in two ways. First, they roughly indicate which relative order between the items in each pair is “more likely” in an optimal solution. The algorithm uses this information to decide whether to perform a full move or a partial move. Second, the simple sum of all these pair-based work functions is a lower bound on the optimal cost, which is useful in analyzing the competitive ratio of Fpm.
Related work.
Better bounds are known for randomized algorithms both in the standard model ($\textsf{LUP}_{\mathrm{S}}$) and the uniform cost model ($\textsf{LUP}_{1}$). For $\textsf{LUP}_{\mathrm{S}}$, a long line of research culminated in a $1.6$-competitive algorithm [20, 25, 2, 6], and a $1.5$-lower bound [27]. The upper bound of $1.6$ is tight in the class of so-called projective algorithms, whose computation is uniquely determined by their behavior on two-item instances [8]. For $\textsf{LUP}_{1}$, the ratio is known to be between $1.5$ [4] and $2.64$ [25].
It is possible to generalize the uniform cost function by distinguishing between the cost of $1$ for following a link during search and the cost of $d\geq 1$ for a swap [25, 4] This model is sometimes called the $P^{d}$ model; in this terminology, our $\textsf{LUP}_{1}$ corresponds to $P^{1}$. While Mtf is $4$-competitive for $P^{1}$, it does not generalize in an obvious way to $d>1$. Other known algorithms for the $P^{d}$ model (randomized and deterministic Counter, RandomReset and TimeStamp) have bounds on competitive ratios that monotonically decrease with growing $d$ [25, 4]. In particular, deterministic Counter achieves ratio $4.56$ when $d$ tends to infinity [7] and for the same setting ($d\to\infty$) a recent result by Albers and Janke [4] shows a randomized algorithm TimeStamp that is $2.24$-competitive.
@@
A variety of List Update variants have been investigated in the literature over the last forty years, including models with lookahead [1], locality of reference [10, 5], parameterized approach [18], algorithms with advice [16], prediction [12], or alternative cost models [19, 23, 24]. A particularly interesting model was proposed recently by Azar et al. [13], where an online algorithm is allowed to postpone serving some requests, but is either required to serve them by a specified deadline or pay a delay penalty.
In summary, List Update is one of canonical problems in the area of competitive analysis, used to experiment with refined models of competitive analysis or to study the effects of additional features. This underscores the need to fully resolve the remaining open questions regarding its basic variants, including the question whether ratio $3$ is attainable for $\textsf{LUP}_{1}$.
@@ 2 Preliminaries
An algorithm has to maintain a list of items, while a sequence $\sigma$ of access requests is presented in an online manner. In each step $t\geq 1$, the algorithm is presented an access request $\sigma^{t}$ to an item in the list. If this item is in a location $\ell$, the algorithm incurs cost $\ell-1$ to access it. (The locations in the list are indexed $1,2,$....) Afterwards, the algorithm may change the list configuration by performing an arbitrary number of swaps of neighboring items, each of cost $1$.
For any algorithm $A$, we denote its cost for processing a sequence $\sigma$ by $A(\sigma)$. The optimal algorithm is denoted by Opt.
@@
For an input $\sigma$ and a pair $\{x,y\}\in\mathcal{P}$, $\sigma_{xy}$ is the subsequence of $\sigma$ restricted to requests to items $x$ and $y$ only. Whenever we say that an algorithm serves input $\sigma_{xy}$, we mean that it has to maintain a list of two items, $x$ and $y$.
@@ 2.1 Work Functions
Figure 1: The evolution of the work function $W^{xy}$. Initially, $W^{xy}$ has its minimum in the state $yx$ and $y\prec x$ (in the list of an online algorithm). Thus, the mode of the pair $\{x,y\}$ is $\alpha$. Next, because of the two requests to $x$, the value of $W^{xy}(yx)$ is incremented, while the value at xyxy remains intact. The mode is thus changed from $\alpha$ to $\beta$ and then to $\gamma$. Finally, when an algorithm moves item $x$ before $y$, the mode of the pair $\{x,y\}$ changes to $\alpha$. The value of $w^{xy}$ (the average of $W^{xy}(xy)$ and $W^{xy}(yx)$) increases by ${\frac{1}{2}}$ whenever the mode changes due to a request.
@@ 2.2 Lower Bound on OPT
2.2 gives us a convenient tool to lower bound $\textsl{{Opt}}(\sigma)$. We define the cost of pair-based Opt in step $t$ as $\sum_{\{x,y\}\in\mathcal{P}}\Delta_{t}w^{xy}$. For a given request sequence $\sigma$, the sum of these costs over all steps is a lower bound on the actual value of $\textsl{{Opt}}(\sigma)$.
@@ 4.3 State Transitions
While we defined 12 possible states (3 modes $\times$ 4 flavors), we will show that $\alpha^{n}$, $\gamma^{d}$, and $\gamma^{o}$ never occur. This clearly holds at the very beginning as all pairs are then in state $\alpha^{d}$.
For succinctness, we also combine some of the remaining states, reducing the number of states to the following six: $\alpha^{d}$, $\beta^{d}$, $\alpha^{oe}$, $\beta^{o}$, $\beta^{ne}$ and $\gamma^{ne}$. (For example, a pair is in state $\alpha^{oe}$ if it is in state $\alpha^{o}$ or $\alpha^{e}$.) In the following we analyze the transitions between them. We start with a simple observation.
@@
The pair $\{z^{*},y\}$ is not swapped due to the request, and thus, by subsection 4.2, its mode transition is $\alpha\to\alpha$, $\beta\to\alpha$, $\gamma\to\beta$.
@@ 4.4 Amortized Analysis
For succinctness, wherever it does not lead to ambiguity, we omit subscripts $t$, i.e., write $\Delta\Phi_{xy}$ and $\Delta w^{xy}$ instead of $\Delta_{t}\Phi_{xy}$ and $\Delta_{t}w^{xy}$. We also omit superscripts $t$ in $\mathcal{P}^{t}_{1}$, $\mathcal{P}^{t}_{2}$, and $\mathcal{P}^{t}_{3}$.
@@
  • $\max\{\beta^{d},\beta^{o}\}\leq\gamma^{ne}+\textstyle{\frac{1}{2}}R$.
@@ 4.4.4 Proof of $R$-competitiveness
Next, assume Fpm performs a full move. Then, $\Delta\textsl{{Fpm}}=[2,2,2,2,2,2]\odot V$. By Table 1,
@@
Recall that Fpm is defined to choose the move that minimizes $\Delta\textsl{{Fpm}}+\sum_{y\prec z^{*}}\Delta\Phi_{z^{*}y}$.
@@
$G_{\mathrm{PM}}$ $=\textstyle{\frac{1}{16}}\cdot\big{[}25-\sqrt{17},\;$ $43-3\sqrt{17},\;$ $1+7\sqrt{17},\;$ $63-7\sqrt{17},\;$ $23+\sqrt{17},\;$ $0\big{]},$
$G_{\mathrm{FM}}$ $=\textstyle{\frac{1}{16}}\cdot\big{[}37+3\sqrt{17},\;$ $27-3\sqrt{17},\;$ $5+3\sqrt{17},\;$ $31-7\sqrt{17},\;$ $23+\sqrt{17},\;$ $0\big{]}.$
@@
$G_{\mathrm{COMB}}=\textstyle{\frac{1}{16}}\cdot\big{[}23+\sqrt{17},\;23+\sqrt{17},\;23+\sqrt{17},\;23+\sqrt{17},\;23+\sqrt{17},\;0\big{]}.$
@@
5 Final Remarks
The most intriguing question left open in our work is whether competitive ratio of $3$ can be achieved. We have shown computationally (see Appendix E) that $3$-competitive algorithms exist for lists with up to $6$ items.
However, even for short lists the definition of such $3$-competitive algorithm remains elusive. For many online problems, the most natural candidate is the generic work function algorithm. This algorithm is $2$-competitive in the $\textsf{LUP}_{\mathrm{S}}$ model [9]. However, our computer-aided calculation of its performance shows that its ratio is larger than $3$ already for $5$ items (see Appendix E). It is $3$-competitive for lists of length up to $4$, though.
We do not know whether the analysis of Fpm is tight. For the specific choice of parameters used in the paper, we verified that Fpm is $3$-competitive for lists of length $3$ (see Appendix C), but not better than $3.04$-competitive for lists of length $5$ (see Appendix D).
The focus of this paper is on the $\textsf{LUP}_{1}$ model (also known as $P^{1}$); we believe that the setting of $d=1$ captures the essence and hardness of the deterministic variant. That said, extending the definition and analysis of Fpm to the $P^{d}$ model (for arbitrary $d$) is an interesting open problem that deserves further investigation.